how do you find the slope & the y-intercept of the equation y= -3x + 2
Answer by stanbon(75887)

how do you find the slope & the y-intercept of the equation y= -3x + 2
---
The y-intercept is the value of "y" when x = 0.
y = -3*0 + 2
y = 0 + 2
y = 2
y-intercept (0,2)
-----------------------
The slope is the amount y changes when x changes by 1.
If x = 0 , y = 2
If x = 1, y = (-3)(1)+2 = -1
---
Y has changed by -3
----
So, slope = -3
====================
Cheers,
Stan H.